# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v1

@@ 1 Introduction
Energy consumption in supercomputers and data centers is the sum of power drawn, among other elements, by computing and storage nodes, the interconnection network, and the cooling system. Considering the most powerful supercomputers, the Exascale barrier was first overrun by the Frontier supercomputer, which remained the first in the TOP500 ranking [35] since June 2022, until El Capitan surpassed it in November 2024. Table 1 contains five of the most powerful supercomputers (in terms of computing power) with energy efficiency information supplied as per the last ranking. Energy efficiency is the ratio between the maximum performance and the average power consumption; the higher it is, the better. As we can see, the highest energy efficiency is achieved by the JUPITER booster machine, which is #4 in the ranking as of June 2025.
@@
The operational costs for supercomputers are concerning because of power availability and sustainability reasons. Green500 [34] ranks TOP500 systems according to their power efficiency rather than the raw performance. If we examine the Green500 ranking for that date, we can see that the aforementioned systems are, in terms of power efficiency, in ranks #13, #42, #68, and #12, respectively. Of course, designing a system with high regard for power efficiency leads to lower operational costs, namely in electricity.
@@
Processors and accelerators used in the most powerful supercomputers have special mechanisms for reducing power consumption when thermal or power constraints are present. Intel processors, for example, introduced the C-states in the Sandy Bridge architecture [26] for advanced power management techniques. CPUs have several states defined, providing lower power consumption at the cost of performance and latency. Operating systems can indicate to the processor to enter a lower power state. However, changing states involves microarchitectural operations and voltage transitions, increasing latencies. During state transition times, CPUs execute no code, affecting performance. Still, even the most energy-saving states consume power. These features aim to keep CPU energy consumption proportional to the usage.
@@
In this context, as computing nodes become more energy proportional, the interconnection network can become a significant fraction of computing system power [1]. Therefore, it is also necessary to design high-performance interconnection networks whose energy consumption is proportional to the traffic they carry and the rest of the system elements (e.g., processors, storage, etc).
@@
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. This would make power consumption more proportional to the network usage. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Fig. 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. Each bin $i$ stores the number of registered inactivity periods that last between $width_{bin}\times i$ and $width_{bin}\times(i+1)$ seconds. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the percentage of the histogram “area” below the bin. Regardless of the distribution on inactivity periods, any power-saving mechanism should take advantage of the longer periods to provide as much energy savings as possible. Moreover, we expect the performance penalty of the mechanism to be as low as possible or within a reasonable margin.
@@
This paper explores the effects of network power management strategies on the performance of interconnection networks. First, we describe the background concepts in network energy reduction (see Section 2). Next, we describe the network power for large Ethernet-based interconnection networks used in supercomputers and data centers, and how it can be applied to BXIv3-based networks (see Section 3). We also show that more efficient interconnection networks are required to achieve energy proportionality. Afterwards, we describe the power model for BXIv3 that we have designed and the proposed energy-saving strategies. This model has been implemented in a powerful simulation framework, allowing large-scale network simulations. Furthermore, we have performed a set of simulation experiments (see Section 4) that show that, with the proposed power model strategies, the network power can be lowered without gravely affecting the overhead on execution time or latency. In particular, we managed to save more than 5 % of energy on the network by focusing on links only, while still outperforming other previous proposals in terms of performance overhead. We manage to improve the PerfBound [28] technique, even to the point of reverting power increases and resulting in actual savings. In our opinion, this is a promising start, given that our proposal and assumptions could be portable to future interconnection network technologies, such BXIv3 and the upcoming (Formerly Intel’s) Omnipath [^1] https://www.theregister.com/2025/06/09/omnipath_is_back/ [8].
@@
    Study and implementation discussion of PerfBound [28] in our network simulator.
@@ 2.1 Energy consumption in high-speed interconnection networks
Network performance is affected by several network design issues, such as the topology, routing, flow control, and operating frequency, among others. Thus, designing an efficient and cost-effective network is crucial to guarantee the performance of the entire supercomputer and datacenter, as well as a balance between this performance and the investment and operational costs. Regarding HPC network technologies, a quick look into the TOP500 list shows us that Gigabit Ethernet, in various revisions, is present in many systems (32.8 %), only second to Infiniband (54.2 %). Such is the momentum of Ethernet in these environments that technologies like iWarp [10] and RoCE (RDMA over Converged Ethernet) [24] are working towards providing Ethernet with RDMA, present in interconnects like Infiniband.
@@
Despite the notable capital investment for the equipment (CAPEX), operation costs (OPEX) and energy constraints are also relevant. The most notable examples of OPEX in HPC are power supply, cooling, and maintenance. Saving energy to reduce OPEX is being tackled by reducing power consumption (at the potential cost of performance), and cooling requirements are proportional to power consumption and thus, heat dissipation. The sheer scale of HPC environments implies huge CAPEX and OPEX. Electricity costs can be a deciding factor for HPC environments. Therefore, it is no surprise that power consumption reduction is sought, especially when examining the sparse network access in scientific applications. In fact, InfiniBand, the most pervasive interconnect in HPC, has power-saving mechanisms [13]. The ubiquity of Ethernet in other commercial areas has also caused several proposals to appear for Ethernet as well [17, 18]. The IEEE Task Force assembled to provide the technology with a common standard, resulting in the standard IEEE 802.3az, also known as Energy Efficient Ethernet (EEE).
@@
Several design and operation strategies have been applied to reduce energy consumption in interconnection networks. Specifically, network energy consumption can be reduced by designing networks with high-radix switches. By increasing the switch radix, fewer routers are required. As switches aggregate more components, this also reduces the number of hops [20]. Consequently, the number of links in the network is reduced when the hop count is lower. Routing algorithms can also be relevant to energy consumption if additional links provide several paths between two nodes. Of course, these links will consume energy when unused.
@@
On the other hand, operation cost (OPEX) in HPC environments increases as more network elements and nodes are added. Like with processors, network energy consumption could be reduced when the network is not being used. Periods of inactivity are determined by the communication pattern of applications. Scientific simulations attempt to minimize inter-node communication, with sparse bursts of network activity. On the other hand, in data centers, we can find deep learning applications, distributed storage, cloud computing, and data-intensive applications (e.g., search engines). The network load for this second set of applications depends on user requests, which need to be served as soon as possible to reduce perceived latency by end users [5].
@@ 2.2 Initial attempts for reducing link power consumption
The power consumption of any device can be divided into static and dynamic power. The later depends on clock frequency, supply voltage and transistor technology. Power consumption is roughly proportional to the cube of the frequency, which is promising for power saving. Dynamic Voltage and Frequency Scaling (DVFS) adjust voltage and frequency based on network utilization to save energy during idle periods and maintain performance under load. Implementations like Adaptive Supply Serial Links (ASSL) adjust supply voltage dynamically, though hardware complexity and additional circuitry such as voltage regulators are required [19, 30]. This conflicts with the objectives of interconnection networks being simple.
Since network technologies have gone through several revisions and can usually work at previous-generation data rates, this was thought to be a good strategy for energy saving. Adaptive Link Rate (ALR) aims to accomplish this by lowering link rates during periods of low utilization without altering voltages or frequencies directly. Instead, links can negotiate lower speeds. Although it can significantly save power, switching link rates introduces latency due to MAC handshakes, PHY configuration, and resynchronization [17], making implementation complex and limiting practical adoption. This is detrimental in HPC environments because it can turn the network into a bottleneck when it is more needed.
@@ 2.3 Low Power Idle (LPI)
As ALR proved inefficient, Low Power Idle (LPI) [18] was proposed. LPI turns components off during inactivity periods instead of modifying the link frequency, which was more suitable. The corresponding states for links are Wake for ports active for transmission, and Sleep for inactive ports in a power-saving state. The underlying idea is that higher data rates provide the most energy-efficient transmission in terms of Joules per bit. Contrary to ALR, the switch between Wake and Sleep states occurs in microseconds instead of nanoseconds. Figure 2 shows these states and how they relate to packet transmissions. Note that ${t}_{w}$ corresponds to the time to switch the port from Sleep to Wake, and ${t}_{s}$ for the opposite transition. Specific values for these transition times on different Ethernet revisions are displayed in Table 2 [32].
@@
The IEEE 802.3az draft [32] specifies values for ${t}_{w}$ and ${t}_{s}$ for several Ethernet specifications, as shown in Table 2. For a comparison, the time to send a 1500-byte frame has also been considered. As we can see, higher speeds reduce frame times, but increase Sleep and Wake transition times considerably. Therefore, LPI is suitable for large bursts followed by large periods of inactivity.
@@
Table 2: Link transition times for different Ethernet versions (extracted from [32]).
@@ 2.4 Energy Efficient Ethernet (EEE)
The IEEE 802.3az standard, also known as Energy Efficient Ethernet (EEE) [11, 32], makes use of the Low Power Idle (LPI) mechanism to reduce energy consumption during idle periods on Ethernet links. This mechanism defines two states: Active and LPI, which involve transitioning between normal operation and a low-power mode. The EEE standard specifies only the basic behavior of LPI, leaving implementation details such as transition policies to the vendors.
When EEE development started, only Sleep and Wake (sometimes referred to as Idle and Active, respectively) were envisioned to provide power savings. Later enhancements to the LPI mechanism were proposed in subsequent amendments to the IEEE 802.3 standard. Specifically, the IEEE 802.3bj amendment, which targets 40 Gb/s and higher-speed Ethernet links, introduces refinements such as the Fast Wake state [7]. Also, the former Sleep state is renamed as Deep Sleep. The Fast Wake state addresses the latency drawbacks of the original LPI proposal by maintaining partial activity in PHY components, while Deep Sleep shuts more components down. This reduces wake transition times significantly, from 4.6 µs in Deep Sleep to around 250-500 ns in Fast Wake. However, this comes at the cost of reduced power savings, typically around 20–40% for Fast Wake compared to up to 90% in Deep Sleep [27]. Anyway, the criterion for designing transitions at the hardware is still left to the vendors.
@@
The different link power levels are a basis for implementing LPI, included in the EEE standard [11]. This has led to several proposals tailored for HPC so that end-to-end latency is not greatly increased by the overhead of turning links on [25, 4]. The idea for HPC environments is to employ Fast Wake so that latency is not greatly affected by ports powering up. However, transitioning up as soon as packet transmission ends causes ports to spend more time transitioning than transmitting packets. Table 3 shows a summary of the available states in EEE, with their link power and transition time.
@@
Table 3: Summary of EEE link power states (based on [7]).
@@ 2.5 Proposals to complement EEE
One option to reduce transitions is frame buffering: accumulating packets on buffers and only transitioning to Wake on certain conditions [11, 27]. Once a packet is on an output port buffer or is granted crossbar traversal on switches, an occupation threshold and a time limit are set. The buffer occupation is set to exploit packet coalescing and thus, avoid transitioning to Wake once per packet. However, a timeout is needed to ensure packets are eventually sent regardless of the buffer occupation. Obviously, this can cause additional latency for every hop. In the worst-case scenario, every communication is composed of small and spaced out packets that do not coincide in time or share network paths. This means that packets will be stalled at every port and have their latency increased both by the buffering timeout and the transition time to Wake. Of course, this is not viable for latency-sensitive applications. As we cannot assume packet coalescing to mitigate the overhead of port transitions, the only option left is to keep links in the Wake state for longer. If the port is left in Wake state for a moderate period after transmission, there is no overhead to a packet that must be transmitted within that time. Power Down Threshold (PDT) [27] was proposed so that every port stays operational for an additional time after transmission, at the cost of potential power saving.
@@
Due to these limitations, ports should establish their own ${t}_{PDT}$ value, based on previous network activity. To solve this issue, Saravanan proposed PerfBound [28]. The basis of PerfBound is forming a histogram on each port with the inactivity periods so that port can calculate its particular ${t}_{PDT}$ at any given moment in a way that degradation is bound. PerfBound also requires every port to store (or calculate) the distance in hops of any packet, given a destination. It uses one counter per possible distance value to calculate the average distance of previous packets. In this way, we can quantify the possible degradation of a packet due to the powering up of multiple ports along its route.

### v2

@@ 1 Introduction
Energy consumption in supercomputers and data centers is the sum of power drawn, among other elements, by computing and storage nodes, the interconnection network, and the cooling system. Considering the most powerful supercomputers, the Exascale barrier was first overrun by the Frontier supercomputer, which remained the first in the TOP500 ranking [top500] since June 2022, until El Capitan surpassed it in November 2024. Table 1 contains five of the most powerful supercomputers (in terms of computing power) with energy efficiency information supplied as per the last ranking. Energy efficiency is the ratio between the maximum performance and the average power consumption; the higher it is, the better. As we can see, the highest energy efficiency is achieved by the JUPITER booster machine, which is #4 in the ranking as of June 2025.
@@
The operational costs for supercomputers are concerning because of power availability and sustainability reasons. Green500 [green500] ranks TOP500 systems according to their power efficiency rather than the raw performance. If we examine the Green500 ranking for that date, we can see that the aforementioned systems are, in terms of power efficiency, in ranks #13, #42, #68, and #12, respectively. Of course, designing a system with high regard for power efficiency leads to lower operational costs, namely in electricity.
@@
Processors and accelerators used in the most powerful supercomputers have special mechanisms for reducing power consumption when thermal or power constraints are present. Intel processors, for example, introduced the C-states in the Sandy Bridge architecture [intelsandybridge] for advanced power management techniques. CPUs have several states defined, providing lower power consumption at the cost of performance and latency. Operating systems can indicate to the processor to enter a lower power state. However, changing states involves microarchitectural operations and voltage transitions, increasing latencies. During state transition times, CPUs execute no code, affecting performance. Still, even the most energy-saving states consume power. These features aim to keep CPU energy consumption proportional to the usage.
@@
In this context, as computing nodes become more energy proportional, the interconnection network can become a significant fraction of computing system power [energyproportional]. Therefore, it is also necessary to design high-performance interconnection networks whose energy consumption is proportional to the traffic they carry and the rest of the system elements (e.g., processors, storage, etc).
@@
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. This would make power consumption more proportional to the network usage. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Figure 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. Each bin $i$ stores the number of registered inactivity periods that last between $width_{bin}\times i$ and $width_{bin}\times(i+1)$ seconds. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the percentage of the histogram “area” below the bin. Regardless of the distribution on inactivity periods, any power-saving mechanism should take advantage of the longer periods to provide as much energy savings as possible. Moreover, we expect the performance penalty of the mechanism to be as low as possible or within a reasonable margin.
@@
This paper explores the effects of network power management strategies on the performance of interconnection networks. First, we describe the background concepts in network energy reduction (see Section 2). Next, we describe the network power for large Ethernet-based interconnection networks used in supercomputers and data centers, and how it can be applied to BXIv3-based networks (see Section 3). We also show that more efficient interconnection networks are required to achieve energy proportionality. Afterwards, we describe the power model for BXIv3 that we have designed and the proposed energy-saving strategies. This model has been implemented in a powerful simulation framework, allowing large-scale network simulations. Furthermore, we have performed a set of simulation experiments (see Section 4) that show that, with the proposed power model strategies, the network power can be lowered without gravely affecting the overhead on execution time or latency. In particular, we managed to save more than 5 % of energy on the network by focusing on links only, while still outperforming other previous proposals in terms of performance overhead. We manage to improve the PerfBound [perfbound] technique, even to the point of reverting power increases and resulting in actual savings. In our opinion, this is a promising start, given that our proposal and assumptions could be portable to future interconnection network technologies, such BXIv3 and the upcoming (Formerly Intel’s) Omnipath [^1] https://www.theregister.com/2025/06/09/omnipath_is_back/ [omnipath].
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. This would make power consumption more proportional to the network usage. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Figure 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the percentage of histogram ”area” below the bin.
@@
    Study and implementation discussion of PerfBound [perfbound] in our network simulator.
@@ 2.1 Energy consumption in high-speed interconnection networks
Network performance is affected by several network design issues, such as the topology, routing, flow control, and operating frequency, among others. Thus, designing an efficient and cost-effective network is crucial to guarantee the performance of the entire supercomputer and datacenter, as well as a balance between this performance and the investment and operational costs. Regarding HPC network technologies, a quick look into the TOP500 list shows us that Gigabit Ethernet, in various revisions, is present in many systems (32.8 %), only second to Infiniband (54.2 %). Such is the momentum of Ethernet in these environments that technologies like iWarp [iwarp] and RoCE (RDMA over Converged Ethernet) [roceiwarp] are working towards providing Ethernet with RDMA, present in interconnects like Infiniband.
@@
Despite the notable capital investment for the equipment (CAPEX), operation costs (OPEX) and energy constraints are also relevant. The most notable examples of OPEX in HPC are power supply, cooling, and maintenance. Saving energy to reduce OPEX is being tackled by reducing power consumption (at the potential cost of performance), and cooling requirements are proportional to power consumption and thus, heat dissipation. The sheer scale of HPC environments implies huge CAPEX and OPEX. Electricity costs can be a deciding factor for HPC environments. Therefore, it is no surprise that power consumption reduction is sought, especially when examining the sparse network access in scientific applications. In fact, InfiniBand, the most pervasive interconnect in HPC, has power-saving mechanisms [dickov_software-managed_2014]. The ubiquity of Ethernet in other commercial areas has also caused several proposals to appear for Ethernet as well [gunaratne_reducing_2008, hays_activeidle_2008]. The IEEE Task Force assembled to provide the technology with a common standard, resulting in the standard IEEE 802.3az, also known as Energy Efficient Ethernet (EEE).
@@
Several design and operation strategies have been applied to reduce energy consumption in interconnection networks. Specifically, network energy consumption can be reduced by designing networks with high-radix switches. By increasing the switch radix, fewer routers are required. As switches aggregate more components, this also reduces the number of hops [kim_microarchitecture_2005]. Consequently, the number of links in the network is reduced when the hop count is lower. Routing algorithms can also be relevant to energy consumption if additional links provide several paths between two nodes. Of course, these links will consume energy when unused.
@@
On the other hand, operation cost (OPEX) in HPC environments increases as more network elements and nodes are added. Like with processors, network energy consumption could be reduced when the network is not being used. Periods of inactivity are determined by the communication pattern of applications. Scientific simulations attempt to minimize inter-node communication, with sparse bursts of network activity. On the other hand, in data centers, we can find deep learning applications, distributed storage, cloud computing, and data-intensive applications (e.g., search engines). The network load for this second set of applications depends on user requests, which need to be served as soon as possible to reduce perceived latency by end users [andujar_extending_2023].
@@ 2.2 Initial attempts for reducing link power consumption
The power consumption of any device can be divided into static and dynamic power. The later depends on clock frequency, supply voltage and transistor technology. Power consumption is roughly proportional to the cube of the frequency, which is promising for power saving. Dynamic Voltage and Frequency Scaling (DVFS) adjust voltage and frequency based on network utilization to save energy during idle periods and maintain performance under load. Implementations like Adaptive Supply Serial Links (ASSL) adjust supply voltage dynamically, though hardware complexity and additional circuitry such as voltage regulators are required [kim_adaptive_2002, shang_dynamic_2003]. This conflicts with the objectives of interconnection networks being simple.
Since network technologies have gone through several revisions and can usually work at previous-generation data rates, this was thought to be a good strategy for energy saving. Adaptive Link Rate (ALR) aims to accomplish this by lowering link rates during periods of low utilization without altering voltages or frequencies directly. Instead, links can negotiate lower speeds. Although it can significantly save power, switching link rates introduces latency due to MAC handshakes, PHY configuration, and resynchronization [gunaratne_reducing_2008], making implementation complex and limiting practical adoption. This is detrimental in HPC environments because it can turn the network into a bottleneck when it is more needed.
@@ 2.3 Low Power Idle (LPI)
As ALR proved inefficient, Low Power Idle (LPI) [hays_activeidle_2008] was proposed. LPI turns components off during inactivity periods instead of modifying the link frequency, which was more suitable. The corresponding states for links are Wake for ports active for transmission, and Sleep for inactive ports in a power-saving state. The underlying idea is that higher data rates provide the most energy-efficient transmission in terms of Joules per bit. Contrary to ALR, the switch between Wake and Sleep states occurs in microseconds instead of nanoseconds. Figure 2 shows these states and how they relate to packet transmissions. Note that ${t}_{w}$ corresponds to the time to switch the port from Sleep to Wake, and ${t}_{s}$ for the opposite transition. Specific values for these transition times on different Ethernet revisions are displayed in Table 2 [specifications_ieee_2009].
@@
The IEEE 802.3az draft [specifications_ieee_2009] specifies values for ${t}_{w}$ and ${t}_{s}$ for several Ethernet specifications, as shown in Table 2. For a comparison, the time to send a 1500-byte frame has also been considered. As we can see, higher speeds reduce frame times, but increase Sleep and Wake transition times considerably. Therefore, LPI is suitable for large bursts followed by large periods of inactivity.
@@
Table 2: Link transition times for different Ethernet versions (extracted from [specifications_ieee_2009]).
@@ 2.4 Energy Efficient Ethernet (EEE)
The IEEE 802.3az standard, also known as Energy Efficient Ethernet (EEE) [christensen_ieee_2010, specifications_ieee_2009], makes use of the Low Power Idle (LPI) mechanism to reduce energy consumption during idle periods on Ethernet links. This mechanism defines two states: Active and LPI, which involve transitioning between normal operation and a low-power mode. The EEE standard specifies only the basic behavior of LPI, leaving implementation details such as transition policies to the vendors.
When EEE development started, only Sleep and Wake (sometimes referred to as Idle and Active, respectively) were envisioned to provide power savings. Later enhancements to the LPI mechanism were proposed in subsequent amendments to the IEEE 802.3 standard. Specifically, the IEEE 802.3bj amendment, which targets 40 Gb/s and higher-speed Ethernet links, introduces refinements such as the Fast Wake state [noauthor_options_2012]. Also, the former Sleep state is renamed as Deep Sleep. The Fast Wake state addresses the latency drawbacks of the original LPI proposal by maintaining partial activity in PHY components, while Deep Sleep shuts more components down. This reduces wake transition times significantly, from 4.6 µs in Deep Sleep to around 250-500 ns in Fast Wake. However, this comes at the cost of reduced power savings, typically around 20–40% for Fast Wake compared to up to 90% in Deep Sleep [saravanan_powerperformance_2013]. Anyway, the criterion for designing transitions at the hardware is still left to the vendors.
@@
The different link power levels are a basis for implementing LPI, included in the EEE standard [christensen_ieee_2010]. This has led to several proposals tailored for HPC so that end-to-end latency is not greatly increased by the overhead of turning links on [reviriego_performance_2009, ANDUJAR2023126]. The idea for HPC environments is to employ Fast Wake so that latency is not greatly affected by ports powering up. However, transitioning up as soon as packet transmission ends causes ports to spend more time transitioning than transmitting packets. Table 3 shows a summary of the available states in EEE, with their link power and transition time.
@@
Table 3: Summary of EEE link power states (based on [noauthor_options_2012]).
@@ 2.5 Proposals to complement EEE
One option to reduce transitions is frame buffering: accumulating packets on buffers and only transitioning to Wake on certain conditions [christensen_ieee_2010, saravanan_powerperformance_2013]. Once a packet is on an output port buffer or is granted crossbar traversal on switches, an occupation threshold and a time limit are set. The buffer occupation is set to exploit packet coalescing and thus, avoid transitioning to Wake once per packet. However, a timeout is needed to ensure packets are eventually sent regardless of the buffer occupation. Obviously, this can cause additional latency for every hop. In the worst-case scenario, every communication is composed of small and spaced out packets that do not coincide in time or share network paths. This means that packets will be stalled at every port and have their latency increased both by the buffering timeout and the transition time to Wake. Of course, this is not viable for latency-sensitive applications. As we cannot assume packet coalescing to mitigate the overhead of port transitions, the only option left is to keep links in the Wake state for longer. If the port is left in Wake state for a moderate period after transmission, there is no overhead to a packet that must be transmitted within that time. Power Down Threshold (PDT) [saravanan_powerperformance_2013] was proposed so that every port stays operational for an additional time after transmission, at the cost of potential power saving.
@@
Due to these limitations, ports should establish their own ${t}_{PDT}$ value, based on previous network activity. To solve this issue, Saravanan proposed PerfBound [perfbound]. The basis of PerfBound is forming a histogram on each port with the inactivity periods so that port can calculate its particular ${t}_{PDT}$ at any given moment in a way that degradation is bound. PerfBound also requires every port to store (or calculate) the distance in hops of any packet, given a destination. It uses one counter per possible distance value to calculate the average distance of previous packets. In this way, we can quantify the possible degradation of a packet due to the powering up of multiple ports along its route.